LII The matrices B and C are given by B= 5 1 ( 22) and C=1 (3 3 . Determine the inverse of B and hence show that CB B:C.
Answer #1

1221 B = T3 3 B1 = determinate of B 2 cans 25) = 24 0 S B is non- singulas. 1.e. Invesse of an exist. B = I adj. B 181 24 (=C 24 | 12 12 / E = L 15-1122 lol s113 3 = - 10-3 24 1- 2415 10-3 - 2 + 15 1 L - Saalis 151 to from o e o est Btc. Scanned_wai